3. A bicycle has a momentum of 25.00 kg* m/s and a velocity of 2.5 m/s . What is the bicycle's
1
Expert's answer
2020-04-26T18:49:31-0400

The mass of the bicycle

"m=\frac{p}{v}=\frac{25 \:\rm kg\cdot m/s}{2.5\:\rm m/s}=10\:\rm kg."